solve the equation 6(d+6)=4(d-6)+2d
Answer by jai_kos(139)

Given 6(d+6)=4(d-6)+2d
 6d + 36 = 4d – 24 + 2d
Get all the “d “ to one side and the numbers to the other side, we get
6d – 4d – 2d = -36 -24
0 = -60
This is inconsistent equation.
Because the terms vanishes.